Circular No. 8126

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 H3 (NEAT) S. H. Pravdo, Jet Propulsion Laboratory, reports the NEAT discovery

on Haleakala images of a comet (discovery observation given below) with a coma diameter of ∼ 14′′ and an unresolved core of diameter ∼ 4′′ or less. Following posting on the NEO Confirmation Page, other observers have also reported the cometary appearance from CCD images, including J. E. McGaha (0.30-m reflector, Tucson, AZ; fainter outer coma of diameter ∼ 10′′ with a brighter core of diameter ∼ 5′′); J. Young (0.6-m reflector, Table Mountain; coma diameter ∼ 8′′, and 16′′ tail in p.a. 250^{o}, affected by cirrus clouds), and P. R. Holvorcem and M. Schwartz (Tenagra IV 0.36-m telescope, near Nogales, AZ; coma diameter 28′′ and m_{1} = 15.4–15.7 on May 1.47).

2003 UT  α_{2000}  δ_{2000}  m_{1}
Apr. 30.45200  19^{h}43^{m}46.^{s}25  −14^{o}40′20′′.7  17.2

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-J02.

T = 2003 Oct. 27.074 TT  ω = 82.^{o}997
Ω = 270.627 }2000.0
q = 1.75468 AU  i = 52.397

V4745 SAGITTARII Red magnitude estimates of V4745 Sgr by W. Liller, Viña del Mar,

Chile, from Technical Pan films taken with an orange filter: Apr. 19.179 UT, 7.8 (thin clouds); 23.197, 8.3 (moon nearby); 27.205, 8.1; 27.208, 8.3.

T. Kato, Kyoto University, reports the following magnitude estimates for V4745 Sgr: Apr. 8.355 UT, V = [14.0 (G. Pojmanski, Warsaw University Observatory; ASAS-3 system, Las Campanas Observatory); 8.790, [12.0 (H. Nishimura, Kakegawa, Shizuoka, Japan; 200-mm f/4 lens + T-Max 400 film); 9.759, [11.0 (Nishimura); 9.80, [12.0 (K. Haseda, Aichi, Japan; Tri-X film); 10.366, V = 10.63 (Pojmanski); 12.338, V = 8.37 (Pojmanski); 12.765, 8.5 (Nishimura); 13.27, 7.3 (R. Y. Shida, São Paulo, Brazil; digital camera); 14.335, V = 8.02 (Pojmanski); 15.353, V = 7.80 (Pojmanski); 18.363, V = 7.41 (Pojmanski); 21.730, 8.5 (Nishimura); 25.708, V = 9.16 (S. Kiyota, Tsukuba, Japan); 26.286, V = 9.26 (Pojmanski); 26.734, 9.2 (Nishimura); 26.762, 9.0 (Haseda); 30.359, V = 9.44 (Pojmanski).

2003 May 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green